OpenStudy (anonymous):

The equation of the tangent line to the curve x^2+y^2=169 at the point (5,-12) is....

OpenStudy (turingtest):

find y' through implicit differentiation can you do that?

OpenStudy (anonymous):

\[2x+2y(dy/dx)=0\]

OpenStudy (anonymous):

then solve for \[dy/dx\]?

OpenStudy (anonymous):

and then dy/dx is the slope right?

OpenStudy (turingtest):

right, sorry for the late reply

OpenStudy (turingtest):

the use that in point-slope form\[y-y_1=\frac{dy}{dx}(x-x_1)\]

OpenStudy (anonymous):

Actually, one can do this problem by a student who has not seen derivative yet.

OpenStudy (anonymous):

The point A=(5,-12) is on the circle x^2+y^2=169. The line AO has slope -12/5 then the slope of the tangent at A is 5/12 The equation of that tangent to the circle at A is y = 5/12(x-5) -12